OpenStudy (anonymous):

a golf ball dropped onto a hard surface rebounds exactly 1/3 of the height from which it falls? how far will it rebound after hitting the surface for the third time if it is intially dropped from a height of 18m?

OpenStudy (anonymous):

well 1st hit = 18 /3 = 6 2nd hit = 6/3 = 2 3rd hit = 2/3 m

OpenStudy (anonymous):

if u wanna do equation this would be y = x / (3n) where y is the final rebound , x is the initial fall height , n is number of surface hits